OpenStudy (anonymous):

A spherical metal dome of radius 15cm is electrically charged. it has a positive charge of 2.5μC distributed uniformly on its surface. Calculate the electric field strength at the surface of the dome.

OpenStudy (stormfire1):

\[E=\frac{Q}{4pi \epsilon_0r^2}\] You know Q and r so you have all the values you need to solve for the field.

OpenStudy (stormfire1):

This may help if you're unclear: http://hyperphysics.phy-astr.gsu.edu/hbase/electric/elesph.html#c3

OpenStudy (anonymous):

E = 2.5 4π8.85*10^ˉ¹² *(0.15)² =2.9*10^10

OpenStudy (anonymous):

Is the calculation correct?

OpenStudy (stormfire1):

Gotta find my calculator...just a minute :)

OpenStudy (stormfire1):

\[\frac{2.5 x 10^{-6}C}{4pi*8.85x10^{-12}*(0.15m)^2}=9.99 x 10^5 N/C\]

OpenStudy (anonymous):

I dont get the same answer i dont know why?

OpenStudy (stormfire1):

Are you typing it in as I did above into your calculator or doing it some other way?

OpenStudy (anonymous):

I,m typing in the same way into the calculator!

OpenStudy (stormfire1):

ok, let me make sure *I* typed into my calculator right...

OpenStudy (anonymous):

Ok

OpenStudy (stormfire1):

Sorry...phone. I still get the same answer

OpenStudy (stormfire1):

2.5E-6/(4*pi*8.85E-12*(0.15)^2) = 999089.4105

OpenStudy (stormfire1):

That's exactly how I typed it in

OpenStudy (anonymous):

Ok i will try it again!

OpenStudy (anonymous):

Yes i get 999089.4105

OpenStudy (stormfire1):

good :) I thought I was going crazy.

OpenStudy (anonymous):

Okay i have a follow on question. Explain how the answer would change at a distance of 30cm from the surface of the dome?

OpenStudy (stormfire1):

Well, if you double the radius in the formula you will get 4 times less field (because the radius is squared in the denominator). You can plug it into the formula to see that.

OpenStudy (anonymous):

I got 249772.3526?

OpenStudy (stormfire1):

Yep... 1/4th of what it was at 15cm

OpenStudy (anonymous):

Ok good. So is the answer 2.5*10^5?

OpenStudy (stormfire1):

Yes

OpenStudy (stormfire1):

I think the question doesn't really want the numerical answer though. I believe it wants you to explain that it will be 4 times less that it was at 15cm because the radius is squared in the denominator.

OpenStudy (stormfire1):

This is a common type of physics question...it wants you to understand how changing the radius (doubling it, halfing it, etc) affects the answer.

OpenStudy (anonymous):

ok thanks